Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
bonnand, gouaty, tacca, eisenmann, flaminio, mwas, masserot, pacaud, letendre, rolland - 23:07 Tuesday 12 March 2019 (45240) Print this report
Installation of HQE photodiodes on SDB2

Today’s shift was dedicated to the installation of the HQE photodiodes on SDB2.

The SDB2 minitower venting started at 09h13 utc and the bench was opened at 09h46 utc.

The photodiode air boxes 09 and 18 were removed from B1-PD2 and B1-PD1 respectively.

The new High Quantum Efficiency photodiodes were installed (air box 07 on B1-PD1 and air box 22 on B1-PD2) and their alignment was checked ok with the single bounce beam from NI. The new photodiodes are equipped with additional beam dumps to dump the reflections from the anti-reflective surfaces of the beam splitter and mirror placed in front of each photodiode.

The optics on B1_M2 and B1_M3 have been swapped in order to restore the configuration with each B1 photodiode receiving 50% of the power. The 50% beam splitter is now installed on B1_M2 and a 99.9% reflective mirror on B1_M3.

The current thresholds for the photodiode shutter closing have been updated for both photodiodes in the SDB2_dbox_bench configuration file. We set the minimum current threshold allowing us to rearm the Vbias on the photodiodes (below these thresholds the Vbias cannot be rearmed) : this corresponds to a threshold equal to 7.25 on B1-PD1 and 8.45 on B1-PD2. Given the offsets in the Ibias current (around 3.35 for PD1 and 4.4 for PD2), these thresholds should correspond to a maximum power equal to 3.9 mA on PD1 and 4.05 mA on PD2.

We also updated the power threshold on B1 for triggering the closing of the SDB1 fast shutter. The new threshold is now 3 mW per photodiode.

We profited from this intervention to check the beam going towards the B1p camera. And we identified that the B1p beam was close to clip on the mount B1p_M1. Thus we displaced the B1p_M1 mount towards north. During this operation, we noticed that the B1p_QD2 quadrant was not responding any more to any request (status, Vbias rearm, shutter open). We noticed a similar issue on all the SDB2 quadrants (B1p and B5). With the help from Alain we could identify the source of this problem as being the network bridge ethmt6 (netcom 413) not responding any more. We thus opened the bench electronic air container in order to investigate the problem, but it was not possible to recover the communication with this ethernet bridge even after some resetting and power on/off attempts. Therefore this part has to be replaced. The replacement is foreseen for tomorrow early afternoon (meanwhile Alain is going to bring us a spare ethernet bridge).

We then let the bench floating again, closed the minitower and started the repumping at about 17h25 utc. The pumping was stopped at about 19h00 utc (Fig.1).

Images attached to this report
Comments to this report:
gouaty - 23:12 Tuesday 12 March 2019 (45241) Print this report

During the intervention on SDB2, we mistuned  the FI waveplate in order to reduce the power sent to the B1 photodiodes. The waveplate was retuned at the end of the activity following the procedure described in entry 44656.

sorrentino - 11:03 Wednesday 13 March 2019 (45245) Print this report

In entry 44656 I forgot to mention the SQZ_B1_CamFar setting during the calibration: gain=0 dB, int time = 50 us. The effect of saturation is not negligible in such conditions, so the calibration is not valid for different settings. This morning I adjusted the waveplate again, see attached plot.

Images attached to this comment
Search
×

Warning

×